# Readers Prefer Outputs of AI Trained on Copyrighted Books over Expert Human Writers

### v3

@@ 2 Discussion
Courts understand this factor to concern the extent to which works produced through copying serve as market substitutes for the original author’s work (?, ?). Our study has shown that AI-generated excerpts from in-context prompted models pre-trained on vast internet corpora (including millions of copyrighted works) were strongly disfavored by expert evaluators for both quality and stylistic fidelity, though lay readers showed no clear preference. By contrast, when models are fine-tuned on curated datasets—consisting solely of individual author’s complete works—both experts and lay readers decisively preferred the AI-generated excerpts over human-written examples. Moreover, the fine-tuned excerpts proved almost undetectable as AI-generated text (particularly when compared to outputs from in-context prompted models), while consistently surpassing human writing in blind evaluations.
@@
The U.S. Copyright Office has recognized that such predicate copying may cause cognizable market harm through competing works that the inputs enable, potentially flooding the market and causing “market dilution” (?). While acknowledging this “market dilution” approach to the fourth fair use factor as “uncharted territory,” the Office determined that both the statutory language and underlying concerns of the Copyright Act warrant this inquiry. “…The speed and scale at which AI systems generate content pose a serious risk of diluting markets for works of the same kind as in their training data. That means more competition for sales of an author’s works and more difficulty for audiences in finding them. If thousands of AI-generated romance novels are put on the market, fewer of the human-authored romance novels that the AI was trained on are likely to be sold…Market harm can also stem from AI models’ generation of material stylistically similar to works in their training data…” The Office emphasized that the effect of the copying impacts extant works by putting them in competition with AI-generated outputs. Copyrighted works become fodder for new productions targeting the same markets. Crucially, the Office did not claim that competing AI outputs copy the inputted works; rather, it examined the economic consequences of the predicate copying that enables these competing outputs. This focus on inputs remains essential because, absent the initial copying, no infringement action exists against flooding markets with independently generated works—if a thousand humans write romance novels after reading Barbara Cartland’s novels, they compete but do not infringe. The Copyright Office’s expansive interpretation of “potential market for or value of the copied work” suggests that fair use might not excuse predicate copying even when it doesn’t show up in the end product, if the copying’s effect substitutes for source works.
@@
In Kadrey v. Meta (?), an infringement action brought by 13 book authors against the copying of their books into the database underpinning Meta’s Llama LLM, Judge Chhabria granted summary judgment to Meta but accepted the theory of market dilution: “[I]ndirect substitution is still substitution: If someone bought a romance novel written by an LLM instead of a romance novel written by a human author, the LLM-generated novel is substituting for the human-written one…This case involves a technology that can generate literally millions of secondary works, with a miniscule fraction of the time and creativity used to create the original works it was trained on. No other use…has anything near the potential to flood the market with competing works the way that LLM training does.” Judge Chhabria effectively provided a road map for what authors would have to show to persuade a court that the AI inputs diluted their markets: First, is the AI system capable of generating such substitutional books now or in the near future? Second, what are the markets for the plaintiffs’ books, and do the AI-generated books compete in those markets? Third, what impact does any such competition have on sales—does it displace those works entirely or merely chip away at demand, and are the effects likely to grow as AI-generated books proliferate and models improve? Fourth, how does the threat to the market for the plaintiffs’ books in a world where LLM developers can copy those books compare to the threat to the market for the plaintiffs’ books in a world where the developers can’t copy them?
@@
These findings suggest that the creation of fine-tuned LLMs consisting of the collected copyrighted works (or a substantial number) of individual authors should not be fair use if the LLM is used to create outputs that emulate the author’s works. As the Copyright Office observed, “[f]ine-tuning…usually narrows down the model’s capabilities and might be more aligned with the original purpose of the copyrighted material,” (?) and thus both less “transformative,” and more likely to substitute for it. By contrast, the LLMs employed for in-context prompting do not target particular authors, and therefore can be put to a great variety of uses that do not risk diluting those authors’ markets. Their claim to fair use seems accordingly stronger. But those models can generate author-emulations, and our study has shown that at least as to lay readers, those outputs can substitute.
@@
A reasonable solution might allow the inclusion of copyrighted works in the general-purpose dataset, but would require the model to implement guardrails that would disable it from generating non-parodic imitations of individual authors’ oeuvres (?, ?, ?).As examples of guardrails, AI developers have implemented “refusal protocols” blocking outputs when prompts request content “in the style of” specific authors. Further, current reinforcement learning techniques can be easily modified to steer models away from stylistic imitation. Another solution, particularly where the lower quality of in-context prompting reduces the prospect of market dilution, might be to condition a ruling of fair use on the prominent disclosure of the output’s AI origin. This solution assumes that the public, informed that the output was not human-authored, will be less inclined to select the AI substitute; transparency should diminish the competition between human-authored and machine-generated offerings. The solution also assumes that in-context prompting will not in the future produce outputs that readers will prefer to human-authored text. Improvements in LLMs (and/or increases in the number of works copied into training data) may come to belie that assumption.
3 Methods
3.1 Experimental Setup
@@
We recruited 28 candidates from top MFA programs in the United States, including from programs such as Iowa Writers Workshop, Helen Zell Writers’ Program at University of Michigan, MFA Program in Creative Writing at New York University, Columbia University School of Arts, and MFA in Creative Writing at BU, paying them \$75 for writing each excerpt. These MFA candidates, along with the LLMs described earlier, emulated the style and voice of 50 award-winning authors representing diverse cultural backgrounds and distinct literary voices (full list in Table 2 in SI). The writing prompt provided to MFA candidates contained three components: (i) 20 sample excerpts spanning an author’s complete body of work, (ii) textual descriptions of the author’s distinctive style and voice, and (iii) detailed content specifications about the original author-written excerpt to be emulated. The selection of the 50 authors and the development of all writing prompts were completed in collaboration with five English Literature PhD students who analyzed each author’s literary voice and created the verbalized style descriptions (prompt shown in Figure 5 in SI).
3.1.1 In-context Prompting
@@ 3.1.2 Fine-tuning
3.2 Human Participants for Writing and Evaluation Task
@@
Our expert sample represents the emerging literary elite whose professional judgment shapes contemporary publishing. Among our recruited MFA candidates, several have since been awarded Stanford Stegner Fellowships (one of the most competitive fiction writing fellowships in the United States), received Rhodes Scholarships, assumed editorial positions at prestigious literary magazines such as Joyland, and secured publishing contracts with major houses including W.W. Norton and Harper Collins. These accomplishments underscore that our experts represent not merely MFA candidates, but established and emerging voices whose aesthetic judgments carry professional authority in the literary marketplace.
All recruited experts were residing in the United States at the time of the study. This geographic restriction was necessary for administrative compliance, as U.S. tax regulations require declaration of compensation above certain thresholds, and the university’s payment infrastructure could not accommodate international participants without Individual Taxpayer Identification Numbers (ITINs) or Social Security Numbers (SSNs). Despite this constraint, our expert sample demonstrated substantial demographic diversity: 65% did not identify as cisgender men, and participants self-identified across multiple ethnic backgrounds, including South Asian, White, East Asian, and Black. This diversity helps ensure that our findings are not artifacts of a narrow demographic perspective.
For the 131 lay readers recruited from Prolific, we restricted participation to native-born residents of English-speaking countries (USA and UK only) ensuring that all evaluators had native-level fluency in English. Additional eligibility criteria included maintaining a 100% task acceptance rate on the platform (indicating high performance on previous tasks) and having a college degree (ensuring a baseline level of literacy). We compensated the lay readers according to Prolific’s wage standards with approximately \$15 per hour for their evaluation time. Unlike experts, lay readers represented general educated audiences without specialized training in creative writing or literary analysis; hence their compensation was lower than experts.
3.3 Blind Evaluation
@@
Beyond selecting a preferred excerpt, both experts and lay readers provided 2–3 sentence explanations grounded in textual evidence to justify their choices  (?) (see Figures 2023 in SI for examples). We required these written rationales for two reasons. First, without observing why a choice was made, the latent factors driving that preference remain unclear—hidden user context, demographic or value variation, task ambiguity, or near-tie similarity can inflate effective label noise. Providing rationales or auxiliary side information has been empirically shown to improve label reliability, data efficiency, and debiasing  (?). Second, for lay readers specifically, written explanations serve as a quality control mechanism, helping us identify participants who may be selecting preferences randomly rather than engaging seriously with the task. Readers who provided generic or off-topic rationales, or whose explanations contradicted their stated preferences, were flagged for potential exclusion.
@@
To ensure annotation quality, we implemented several attention checks. We recorded timestamps for each evaluation to identify and exclude participants who rushed through tasks (spending less than 30 seconds per comparison). Additionally, we screened all written rationales using Pangram, [^5] https://www.pangram.com/ a state-of-the-art AI detection tool, and excluded any participants who used generative AI to compose their responses  (?). This screening was necessary because AI-generated rationales would undermine the validity of human preference judgments. Our study was approved by the University of Michigan IRB (HUM00264127) and was preregistered at OSF. [^6] https://osf.io/zt4ad Informed consent was obtained from all participants. As noted earlier, we compensated experts \$75 per 10-excerpt batch for quality judgments and \$100 per 10-excerpt batch for stylistic fidelity evaluations (the higher rate reflecting the additional cognitive load of comparing against original author excerpts).
3.4 Empirical Analyses
@@
We tested three preregistered hypotheses using logistic regression models with CR2 cluster-robust standard errors clustered by reader. Clustering accounts for the fact that each reader evaluated multiple excerpt pairs, which induces within-reader correlation in preference judgments. The CR2 adjustment provides more accurate inference in settings with a modest number of clusters (28 expert readers, 131 lay readers)  (?).
@@
For H1 (can baseline LLMs match human expert performance?), we compared human writing to the average performance across GPT-4o, Claude 3.5 Sonnet, and Gemini 1.5 Pro in the in-context prompting condition. Because each human excerpt was paired with only one AI excerpt from a single model, we pooled AI excerpts across the three models and tested whether readers showed systematic preference for human versus AI writing on average. For H2 (can fine-tuned models match human expert performance?), we directly compared fine-tuned GPT-4o excerpts to human writing. Contrasts were computed separately for expert and lay readers within each outcome (stylistic fidelity, writing quality), yielding four tests per hypothesis (2 reader groups $\times$ 2 outcomes). To control the family-wise error rate across these multiple comparisons, we applied Holm correction separately within each hypothesis-outcome combination. This procedure adjusts $p$-values to maintain valid inference when testing multiple contrasts that address the same research question.
@@
Anonymized trial-level data supporting the findings of this study will be made publicly available on paper acceptance. If you need them for research purpose, please email the first author. Additional details are provided in SI, Section S10.
@@
All analysis and figure-generation code will be made publicly available on paper acceptance. Core analyses can be reproduced by running the numbered R scripts in sequence. Analyses were conducted in R 4.3.1 with key packages including clubSandwich and emmeans. Exact package versions and run instructions are provided in SI, Section S10.
@@ 16 Haruki Murakami (✓) 34 Orhan Pamuk (✓)
18 Ian McEwan (✓) 36 Percival (✓)Everett
@@
S1.3 Finetuning details
Figure 6: The pipeline used to finetune ChatGPT on an author’s entire oeuvre
For finetuning we bought digital ePub versions of these authors’ books and transformed them into plain text files. If the epub file was basically a wrapper around scanned page images, then we ignored that epub. The number of books written by each author vary a lot. For instance Tony Tulathimutte has written only two books Private Citizens and Rejection so we could only finetune GPT-4o on two of them. While for Haruki Murakami we could finetune on 22 books A Wild Sheep Chase, After Dark, After the Quake, Blind Willow, Sleeping Woman, Colorless Tsukuru Tazaki and his Years of Pilgrimage, Dance Dance Dance, First Person Singular, Hard-boiled Wonderland and the End of the World, Hear the Wind Sing, Kafka on the Shore, Killing Commendatore, Men Without Women, Norwegian Wood, Novelist as a Vocation, One and Two, Pinball, 1973, South of the Border, West of the Sun, Sputnik Sweetheart, The City and Its Uncertain Walls, The Elephant Vanishes, The Wind-Up Bird Chronicle, What I Talk About When I Talk About Running, Wind/Pinball: Two Novels
Our entire finetuning pipeline can be seen in Figure 6. We first convert the epub files to txt using https://github.com/kevinboone/epub2txt2. We then segment the entire book into context independent excerpts. At a first pass we naively split the entire book text by existing double-newlines and rejoin them to enforce excerpt size bounds(250-650 words). For the rare cases where the naive splitting would lead to excerpts longer than 650 words we used GPT-4o to segment them again using the prompt Segment it into excerpts of minimum length 300-350 words such that each excerpt is grammatical from the start and doesn’t feel abruptly cut off. There should be zero deletion and break into excerpts at grammatically natural places. Maintain the original word count. Avoid breaking into too many small excerpts. Start directly. Don’t say Here’s or Here is ….
@@
After finetuning is completed at inference time we can then simply generate an excerpt conditioned on a similar instruction that contains a novel content. We specifically excluded the original author written excerpt used for our experiments during fine-tuning so that the model does not get an unfair advantage. We also ensured that the generated outputs do not contain any memorized snippets from the original author written excerpt. In the rare occasion that a model regurgitated verbatim snippets/ngrams from original author written text we resampled it and manually verified that there is no verbatim overlap before using it for evaluation. While supervised fine-tuning in most cases ensures that the generated output contains all the information in the extracted content details, in the rare occasion that it does not we resample/regenerate it again. Last but not least sometimes supervised fine-tuning can lead to ungrammatical output or output with minor inconsistencies. To make sure these mistakes don’t impact human evaluation, we performed a post processing step using GPT-4o using the following prompt Fix grammar, tense, typo, spelling or punctuation error or any other awkward construction/ logical inconsistency
@@
For Margaret Atwood, Kazuo Ishiguro, Salman Rushdie and Haruki Murakami we finetuned GPT-4o for 1 epoch as they had multiple books/longer books. For the rest of the authors we finetuned for 3 epochs. We use default finetuning parameters set by OpenAI in addition to setting batch size as 1 and LR multiplier as 2. Table 19 shows that finetune models don’t regurgitate a lot of verbatim text from the training corpus (author’s entire oeuvre). We should also note that some of these words are also part of the Content in the instruction which would anyway appear in the text and should not be penalized. For all generation(prompting or fine-tuning default temperature=1.0 was used)
@@
Figure 12: Evaluation Interface for writing quality evaluation where readers (expert and lay) just see two excerpts without any disclaimer of its source and choose their preference supported by a reason
Figures 12 and 13 shows the evaluation interface shown to both lay and expert readers. Readers read two excerpts for quality evaluation and three excerpts for stylistic fidelity evaluation given that the stylistic emulation is evaluated with respect to the original author written excerpt.
Figure 13: Evaluation Interface for stylistic fidelity evaluation where readers (expert and lay) just see three excerpts (original and two emulations) without any disclaimer of its source and choose their preference supported by a reason
@@
As can be seen in 14 writing quality has a strong correlation with a faithful style/voice fidelity for In-context Prompting condition compared to Finetuning. The high expert correlation for In-context Prompting suggests that models cannot emulate an authors style properly just by prompting and that experts use stylistic cues as a major factor in quality judgments - likely detecting ”AI-ness” (clichés, purple prose, too much exposition, lack of subtext, mixed metaphors) in the style that makes them rate quality lower too. With finetuning, experts can appreciate quality independent of stylistic fidelity, suggesting the as AI loses its tell-tale signs experts can evaluate each dimension on its own merits. For lay readers the correlation only drops slightly from 0.52 (In-context prompting) to 0.45 (finetuned), compared to experts’ dramatic drop from 0.64 to 0.39. This also mean that lay readers might focus on more surface-level features such as grammar, clarity or flow rather than subtle stylistic markers that distinguishes AI from human writing.

### v4

@@ Discussion
Courts understand this factor to concern the extent to which works produced through copying serve as market substitutes for the original author’s work (?, ?).
Our study has shown that AI-generated excerpts from in-context prompted models pre-trained on vast internet corpora (including millions of copyrighted works) were strongly disfavored by MFA-trained readers for both quality and stylistic fidelity, while college-educated general readers showed no preference on stylistic fidelity but already favored AI-generated text for writing quality. When models are fine-tuned on curated datasets consisting solely of individual authors’ complete works, both groups decisively preferred the AI-generated excerpts over human-written examples. Notably, college-educated general readers, who are generally representative of the book-purchasing public, preferred fine-tuned AI by a substantially wider margin than MFA-trained readers (e.g., writing quality OR = 5.42 vs. 1.87). Moreover, the fine-tuned excerpts proved almost undetectable as AI-generated text (particularly when compared to outputs from in-context prompted models), while both MFA-trained and college-educated general readers preferred them over human writing in blind evaluations.
@@
The U.S. Copyright Office has recognized that such predicate copying may cause cognizable market harm through competing works that the inputs enable, potentially flooding the market and causing “market dilution” (?). While acknowledging this “market dilution” approach to the fourth fair use factor as “uncharted territory,” the Office determined that both the statutory language and underlying concerns of the Copyright Act warrant this inquiry. “…The speed and scale at which AI systems generate content pose a serious risk of diluting markets for works of the same kind as in their training data. That means more competition for sales of an author’s works and more difficulty for audiences in finding them. If thousands of AI-generated romance novels are put on the market, fewer of the human-authored romance novels that the AI was trained on are likely to be sold…Market harm can also stem from AI models’ generation of material stylistically similar to works in their training data…” The Office emphasized that the effect of the copying impacts extant works by putting them in competition with AI-generated outputs. Copyrighted works become fodder for new productions targeting the same markets (?). Crucially, the Office did not claim that competing AI outputs copy the inputted works; rather, it examined the economic consequences of the predicate copying that enables these competing outputs. This focus on inputs remains essential because, absent the initial copying, no infringement action exists against flooding markets with independently generated works—if a thousand humans write romance novels after reading Barbara Cartland’s novels, they compete but do not infringe. The Copyright Office’s expansive interpretation of “potential market for or value of the copied work” suggests that fair use might not excuse predicate copying even when it doesn’t show up in the end product, if the copying’s effect substitutes for source works.
@@
In Kadrey v. Meta (?), an infringement action brought by 13 book authors against the copying of their books into the database underpinning Meta’s Llama LLM, Judge Chhabria granted summary judgment to Meta but accepted the theory of market dilution: “[I]ndirect substitution is still substitution: If someone bought a romance novel written by an LLM instead of a romance novel written by a human author, the LLM-generated novel is substituting for the human-written one…This case involves a technology that can generate literally millions of secondary works, with a miniscule fraction of the time and creativity used to create the original works it was trained on. No other use…has anything near the potential to flood the market with competing works the way that LLM training does.” Judge Chhabria effectively provided a road map for what authors would have to show to persuade a court that the AI inputs diluted their markets: First, is the AI system capable of generating such substitutional books now or in the near future? Second, what are the markets for the plaintiffs’ books, and do the AI-generated books compete in those markets? Third, what impact does any such competition have on sales—does it displace those works entirely or merely chip away at demand, and are the effects likely to grow as AI-generated books proliferate and models improve? Fourth, how does the threat to the market for the plaintiffs’ books differ between a world in which LLM developers may copy those books and a world in which they may not?
@@
These findings suggest that the creation of fine-tuned LLMs consisting of the collected copyrighted works (or a substantial number) of individual authors should not be fair use if the LLM is used to create outputs that emulate the author’s works (?). As the Copyright Office observed, “[f]ine-tuning…usually narrows down the model’s capabilities and might be more aligned with the original purpose of the copyrighted material,” (?) and thus both less “transformative,” and more likely to substitute for it. By contrast, the LLMs employed for in-context prompting do not target particular authors, and therefore can be put to a great variety of uses that do not risk diluting those authors’ markets. Their claim to fair use seems accordingly stronger. But those models can generate author-emulations, and our study has shown that among college-educated general readers, who more closely represent the book-purchasing public, both in-context quality outputs and fine-tuned outputs show clear AI preference, indicating meaningful substitution potential even without author-specific training. Our findings have also been independently corroborated by a large-scale blind quiz conducted by The New York Times in which more than 86,000 readers compared AI and human-written passages and 54% preferred AI-generated text (?), as well as independent reporting in The New Yorker where even accomplished novelists struggled to distinguish fine-tuned AI emulations from human writing (?).
@@
A reasonable solution might allow the inclusion of copyrighted works in the general-purpose dataset, but would require the model to implement guardrails that would disable it from generating non-parodic imitations of individual authors’ oeuvres (?, ?, ?). As examples of guardrails, AI developers have implemented “refusal protocols” blocking outputs when prompts request content “in the style of” specific authors. Further, current reinforcement learning techniques can be easily modified to steer models away from stylistic imitation. However, such guardrails remain imperfect: refusal protocols can be circumvented through prompt rephrasing, and reinforcement learning–based steering may not generalize across all forms of stylistic emulation. Another solution, particularly where the lower quality of in-context prompting reduces the prospect of market dilution, might be to condition a ruling of fair use on the prominent disclosure of the output’s AI origin. This solution assumes that the public, informed that the output was not human-authored, will be less inclined to select the AI substitute; transparency should diminish the competition between human-authored and machine-generated offerings.
Whether disclosure would in fact diminish the competitive advantage of AI-generated works is a question our study cannot directly address, because all evaluations were conducted under blinded conditions. Whether the strong preferences observed here would persist when readers know the text is AI-generated remains an open empirical question requiring a separate, unblinded experimental design. Moreover, even if disclosure were to reduce the quality preference to some degree, relative cost may independently sustain consumer substitution. If AI-generated works enter the market at substantially lower price points than human-authored works—as our cost analysis suggests is feasible—consumers may select the AI version even knowing its provenance, because the quality-to-price ratio remains attractive. The interaction of disclosure, price, and reader preference in actual market conditions therefore warrants further investigation.
Methods
Experimental Setup
@@ Methods
We recruited 28 expert writers affiliated with top MFA programs in the United States, including from programs such as Iowa Writers Workshop, Helen Zell Writers’ Program at University of Michigan, MFA Program in Creative Writing at New York University, Columbia University School of Arts, and MFA in Creative Writing at BU, paying them \$75 for writing each excerpt. These MFA candidates, along with the LLMs described earlier, emulated the style and voice of 50 award-winning authors representing diverse cultural backgrounds and distinct literary voices (full list in Table 2 in SI). The writing prompt provided to MFA candidates contained three components: (i) 20 sample excerpts spanning an author’s complete body of work, (ii) textual descriptions of the author’s distinctive style and voice, and (iii) detailed content specifications about the original author-written excerpt to be emulated. The selection of the 50 authors and the development of all writing prompts were completed in collaboration with five English Literature PhD students who analyzed each author’s literary voice and created the verbalized style descriptions in addition to carefully choosing representative excerpts written by the original author (prompt shown in Figure 5 in SI). For the purpose of the study, MFA-trained expert writers had no time limitation to submit these excerpts and were encouraged to take as long as required to produce their best work.
In-context Prompting
@@
Human Participants for Writing and Evaluation Task
@@
Our MFA-trained writer sample represents the emerging literary elite whose professional judgment shapes contemporary publishing. Among our recruited MFA candidates, several have since been awarded Stanford Stegner Fellowships (one of the most competitive fiction writing fellowships in the United States), received Rhodes Scholarships, assumed editorial positions at prestigious literary magazines such as Joyland, been awarded the Publishers Weekly Star Watch 2025 Honor and secured publishing contracts with major houses including W.W. Norton and Harper Collins. These accomplishments underscore that our experts represent not merely MFA candidates, but established and emerging voices whose aesthetic judgments carry professional authority in the literary marketplace.
All recruited MFAs were residing in the United States at the time of the study. This geographic restriction was necessary for administrative compliance, as U.S. tax regulations require declaration of compensation above certain thresholds, and the university’s payment infrastructure could not accommodate international participants without Individual Taxpayer Identification Numbers (ITINs) or Social Security Numbers (SSNs). Despite this constraint, our expert sample demonstrated substantial demographic diversity: 65% did not identify as cisgender men, and participants self-identified across multiple ethnic backgrounds, including South Asian, White, East Asian, and Black. This diversity helps ensure that our findings are not artifacts of a narrow demographic perspective.
For the 516 college-educated general readers recruited from Prolific, we restricted participation to native-born residents of English-speaking countries (USA and UK only) ensuring that all evaluators had native-level fluency in English. Additional eligibility criteria included maintaining a 100% task acceptance rate on the platform (indicating high performance on previous tasks) and having a college degree (ensuring a baseline level of literacy). We compensated the college-educated general readers according to Prolific’s wage standards with approximately \$15 per hour for their evaluation time. Unlike MFA-trained readers, college-educated general readers represented general educated audiences without specialized training in creative writing or literary analysis; hence their compensation was lower than MFAs.
Blind Evaluation
@@
Beyond selecting a preferred excerpt, both groups of readers provided 2–3 sentence explanations grounded in textual evidence to justify their choices  (?) (see Figures 2023 in SI for examples). We required these written rationales for two reasons. First, without observing why a choice was made, the latent factors driving that preference remain unclear—hidden user context, demographic or value variation, task ambiguity, or near-tie similarity can inflate effective label noise. Providing rationales or auxiliary side information has been empirically shown to improve label reliability, data efficiency, and debiasing  (?). Second, for college-educated general readers specifically, written explanations serve as a quality control mechanism, helping us identify participants who may be selecting preferences randomly rather than engaging seriously with the task. Readers who provided generic or off-topic rationales, or whose explanations contradicted their stated preferences, were flagged for potential exclusion.
@@
To ensure annotation quality, we implemented several attention checks. We recorded timestamps for each evaluation to identify and exclude participants who rushed through tasks (spending less than 30 seconds per comparison). Additionally, we screened all written rationales using Pangram, [^5] https://www.pangram.com/ a state-of-the-art AI detection tool, and excluded any participants who used generative AI to compose their responses  (?). This screening was necessary because AI-generated rationales would undermine the validity of human preference judgments. Our study was approved by the University of Michigan IRB (HUM00264127) and was preregistered at OSF. [^6] https://osf.io/zt4ad Informed consent was obtained from all participants. As noted earlier, we compensated MFAs \$75 per 10-excerpt batch for quality judgments and \$100 per 10-excerpt batch for stylistic fidelity evaluations (the higher rate reflecting the additional cognitive load of comparing against original author excerpts).
Empirical Analyses
@@
We tested three preregistered hypotheses using logistic regression models with CR2 cluster-robust standard errors clustered by reader. Clustering accounts for the fact that each reader evaluated multiple excerpt pairs, which induces within-reader correlation in preference judgments. The CR2 adjustment provides more accurate inference in settings with a modest number of clusters (28 MFA-trained expert readers, 516 college-educated general readers)  (?).
@@
For H1 (can baseline LLMs match human expert performance?), we compared human writing to the average performance across GPT-4o, Claude 3.5 Sonnet, and Gemini 1.5 Pro in the in-context prompting condition. Because each human excerpt was paired with only one AI excerpt from a single model, we pooled AI excerpts across the three models and tested whether readers showed systematic preference for human versus AI writing on average. For H2 (can fine-tuned models match human expert performance?), we directly compared fine-tuned GPT-4o excerpts to human writing. Contrasts were computed separately for MFA-trained readers and college-educated general readers within each outcome (stylistic fidelity, writing quality), yielding four tests per hypothesis (2 reader groups $\times$ 2 outcomes). To control the family-wise error rate across these multiple comparisons, we applied Holm correction separately within each hypothesis-outcome combination. This procedure adjusts $p$-values to maintain valid inference when testing multiple contrasts that address the same research question.
@@
Anonymized trial-level data supporting the findings of this study are available at GitHub at the link (https://github.com/tuhinjubcse/Author-Style-Personalization/tree/main). Additional details are provided in SI, Section S10.
@@
All analysis and figure-generation code is available at GitHub (https://github.com/tuhinjubcse/Author-Style-Personalization/tree/main). Core analyses can be reproduced by running the numbered R scripts in sequence. Analyses were conducted in R 4.3.1 with key packages including clubSandwich and emmeans. Exact package versions and run instructions are provided in SI, Section S10.
@@ 16 Haruki Murakami (✓) 34 Orhan Pamuk (✓)
18 Ian McEwan (✓) 36 Percival Everett (✓)
@@
S1.3 Fine-tuning details
Figure 6: The pipeline used to fine-tune ChatGPT on an author’s entire oeuvre
For fine-tuning we bought digital ePub versions of these authors’ books and transformed them into plain text files. If the epub file was basically a wrapper around scanned page images, then we ignored that epub. The number of books written by each author vary a lot. For instance Tony Tulathimutte has written only two books Private Citizens and Rejection so we could only fine-tune GPT-4o on two of them. While for Haruki Murakami we could fine-tune on 22 books A Wild Sheep Chase, After Dark, After the Quake, Blind Willow, Sleeping Woman, Colorless Tsukuru Tazaki and his Years of Pilgrimage, Dance Dance Dance, First Person Singular, Hard-boiled Wonderland and the End of the World, Hear the Wind Sing, Kafka on the Shore, Killing Commendatore, Men Without Women, Norwegian Wood, Novelist as a Vocation, One and Two, Pinball, 1973, South of the Border, West of the Sun, Sputnik Sweetheart, The City and Its Uncertain Walls, The Elephant Vanishes, The Wind-Up Bird Chronicle, What I Talk About When I Talk About Running, Wind/Pinball: Two Novels
Our entire fine-tuning pipeline can be seen in Figure 6. We first convert the epub files to txt using https://github.com/kevinboone/epub2txt2. We then segment the entire book into context independent excerpts. At a first pass we naively split the entire book text by existing double-newlines and rejoin them to enforce excerpt size bounds(250-650 words). For the rare cases where the naive splitting would lead to excerpts longer than 650 words we used GPT-4o to segment them again using the prompt Segment it into excerpts of minimum length 300-350 words such that each excerpt is grammatical from the start and doesn’t feel abruptly cut off. There should be zero deletion and break into excerpts at grammatically natural places. Maintain the original word count. Avoid breaking into too many small excerpts. Start directly. Don’t say Here’s or Here is ….
@@
After fine-tuning is completed at inference time we can then simply generate an excerpt conditioned on a similar instruction that contains a novel content. We specifically excluded the original author written excerpt used for our experiments during fine-tuning so that the model does not get an unfair advantage. We also ensured that the generated outputs do not contain any memorized snippets from the original author written excerpt. In the rare occasion that a model regurgitated verbatim snippets/ngrams from original author written text we resampled it and manually verified that there is no verbatim overlap before using it for evaluation. While supervised fine-tuning in most cases ensures that the generated output contains all the information in the extracted content details, in the rare occasion that it does not we resample/regenerate it again. Last but not least sometimes supervised fine-tuning can lead to ungrammatical output or output with minor inconsistencies. To make sure these mistakes don’t impact human evaluation, we performed a post processing step using GPT-4o using the following prompt Fix grammar, tense, typo, spelling or punctuation error or any other awkward construction/ logical inconsistency
@@
For Margaret Atwood, Kazuo Ishiguro, Salman Rushdie and Haruki Murakami we fine-tuned GPT-4o for 1 epoch as they had multiple books/longer books. For the rest of the authors we fine-tuned for 3 epochs. We use default fine-tuning parameters set by OpenAI in addition to setting batch size as 1 and LR multiplier as 2. Figure 19 shows that fine-tuned models don’t regurgitate a lot of verbatim text from the training corpus (author’s entire oeuvre). We should also note that some of these words are also part of the Content in the instruction which would anyway appear in the text and should not be penalized. For all generation(prompting or fine-tuning default temperature=1.0 was used)
@@
Figure 12: Evaluation Interface for writing quality evaluation where readers just see two excerpts without any disclaimer of its source and choose their preference supported by a reason
Figures 12 and 13 shows the evaluation interface shown to both MFA-trained and college-educated general readers. Readers read two excerpts for quality evaluation and three excerpts for stylistic fidelity evaluation given that the stylistic emulation is evaluated with respect to the original author written excerpt.
Figure 13: Evaluation Interface for stylistic fidelity evaluation where readers just see three excerpts (original and two emulations) without any disclaimer of its source and choose their preference supported by a reason
@@
As can be seen in Figure 14 writing quality has a strong correlation with a faithful style/voice fidelity for In-context Prompting condition compared to Fine-tuning. The high expert correlation for In-context Prompting suggests that models cannot emulate an author’s style properly just by prompting and that experts use stylistic cues as a major factor in quality judgments - likely detecting “AI-ness” (clichés, purple prose, too much exposition, lack of subtext, mixed metaphors) in the style that makes them rate quality lower too. With fine-tuning, experts can appreciate quality independent of stylistic fidelity, suggesting that as AI loses its tell-tale signs experts can evaluate each dimension on its own merits. For college-educated general readers the correlation only drops slightly from 0.52 (In-context prompting) to 0.45 (fine-tuned), compared to experts’ dramatic drop from 0.64 to 0.39. This also means that college-educated general readers might focus on more surface-level features such as grammar, clarity or flow rather than subtle stylistic markers that distinguish AI from human writing.